OpenStudy (anonymous):

Can anyone design me a cyclic function which repeats only every 360 degrees and has defined values for all angles using normal trig functions?

OpenStudy (anonymous):

Is this possible using real numbers? I doubt it very much, but it can never hurt to try (I have no idea whether using complex numbers would help.

OpenStudy (anonymous):

sin, cos...?

OpenStudy (anonymous):

No, because they are symmetrical, so sin(180-x)=sinx, and I'm seeking a function with ONLY ???x=???x+360

OpenStudy (anonymous):

 like that?

OpenStudy (anonymous):

\[\cos ( \theta \div 2 )\]

OpenStudy (anonymous):

No, you see this thing cosx=cos360-x also, not just + x (this problem exists no matter how you scale it, in this case it would be 720 or 180)

OpenStudy (anonymous):

sin x= sin (360+x) ??